Hello
I need some assistance with this one Ta. It is asking me to: Write a full solution for the following:
A journey takes 1.5 hours longer at 50kph than it does at 75kph. Find the length of the journey.
Thank you
Dave

Answer by checkley75(3666)

50(t+1.5)=75t
50t+75=75t
75t-50t=75
25t=75
t=75/25
t=3 length of time for the trip.
75*3=225 k. is the length of the trip
proof
50*4.5=75*3
225=225